OpenStudy (anonymous):

Indicate which property is illustrated in Step 1. Step 1 7 + (2 + 11) + 0 = 7 + (11 + 2) + 0 Step 2 = (7 + 11) + (2 + 0) Step 3 = (7 + 11) + 2 Step 4 = 18 + 2 A. associative property of addition B. identity property of addition C. distributive property D. commutative property of addition

OpenStudy (anonymous):

@Opcode

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

Indicate which property is illustrated in Step 2. Step 1 2 + (3 + 8) + 0 = 2 + (8 + 3) + 0 Step 2 = (2 + 8) + (3 + 0) Step 3 = (2 + 8) + 3 Step 4 = 10 + 3 A. associative property of addition B. identity property of addition C. commutative property of multiplication D. associative property of multiplication

OpenStudy (anonymous):

@Opcode